Find the x-intercept of the function f(x)=a^x+1 - 1
A. (0,0)
B. (a, 0)
C. (-1, 0)
D. (a-1,0)

Found 2 solutions by rapaljer, stanbon:
Answer by rapaljer(4671)

The x-intercept is the value where f(x)=0.

0=a^(x+1) -1
1= a^(x+1)

Since a^0=1, the exponent must be 0. Therefore
x+1=0
x=-1

The x-intercept is (-1,0).

Happy New Year!! My Email address is rapaljer@seminolestate.edu if you need to contact me!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus


Answer by stanbon(75887)

Find the x-intercept of the function f(x)=a^(x+1)- 1
------
Let y = 0 and solve for "x":
a^(x+1) = 1
x+1 = 0
x = -1
------
Answer: C
=================
-
A. (0,0)
B. (a, 0)
C. (-1, 0)
D. (a-1,0)